Grade 12th passPhysical Chemistry

400 ml of oxygen at 27 c are cooled to -15c at constant pressure .calculate the contraction

m s jadoun
11 Years agoGrade 12th pass

1 Answer

Chirag Goel
11 Years ago
V`/V =   T`/T = 288/300                                                                                                                                                   V` =  (400 x 288)/300 = 384 ml.                                                                                                                                       contraction = 16 ml = 16/400 x 100 = 4`/.  .